# Topology of shallow-water waves on the rotating sphere

### v1

@@ Abstract
${}^{1}$niv Lyon, ENS de Lyon, Univ Claude Bernard, CNRS, Centre de Recherche Astrophysique de Lyon (UMR CNRS 5574), F-69230 Saint-Genis-Laval, France
${}^{2}$niv Lyon, ENS de Lyon, Univ Claude Bernard, CNRS, Laboratoire de Physique (UMR CNRS 5672), F-69342 Lyon, France
@@ 1 Introduction
After noticing the strong resemblance of this $+2$ spectral flow of modes with the topologically-protected modes crossing the gap in certain insulating materials, [1] applied the same arguments used in condensed matter physics to define topological invariants associated to this spectral flow, the Chern numbers or topological charges. These invariants are similar to the ones derived in [14] in the context of the quantum Hall effect. They characterise the singularities in the winding of the eigenvectors of Kelvin’s solutions, which are computed for constant $f$ and thus require only to diagonalise a 3-by-3 matrix. Nevertheless, they happen to be equal to the number of modes gained by the equivalent wavebands in the more elaborated $\beta$-plane model. This counter-intuitive result, which connects the topological properties of Kelvin’s $f$-plane modes to a spectral property of Matsuno’s $\beta$-plane spectrum, is a consequence of the index theorem [15, 16, 17]. It has been applied to predict the conditions of existence and number of modes transiting across the frequency gap in a variety of geophysical and astrophysical models [18, 19, 20, 21], and other physical systems from many different domains such as plasma physics [22, 23], photonics [24], active matter [25, 26] and mechanical metamaterials [27, 28], among others.
However, the equatorial $\beta$-plane approximation used in [13, 1] (and many other studies of geophysical fluid dynamics) is limited in the sense that it only accounts for the curvature of the surface in the variation of the Coriolis parameter, at linear order, and not in the metric. It thus ignores the existence of the poles, and generates a spectrum of solutions defined on the unbounded domain across the equator, which seems paradoxical. Nevertheless, the solutions are accurate as long as they remain trapped around the equator, in the so-called Yoshida waveguide [29, 13], where the linear approximation holds. The trapping length, called the equatorial radius of deformation, is given by
@@
where $c$ is the speed of surface waves without rotation. $L_{\rm eq}$ thus needs to be much smaller than $R$, the radius of the planet. This condition is satisfied for fast-rotating planets, compared to the timescale of propagation of the waves at their surface. For the first baroclinic mode (i.e. the fastest one) in the equatorial ocean on Earth, the equatorial radius of deformation is approximately $300$ kilometers (see e.g. [3], p. 304), so the $\beta$-plane approximation is quite accurate in this context. However, when it comes to global oscillations of larger scales, both the curved metric and the actual sine variation of $f$ with latitude must be taken into account. Moreover, the quantisation of the zonal wave number $k$ cannot be ignored, especially when the wavelength is not small compared to the radius $R$. The full spectrum of shallow-water waves on the sphere is a peculiar problem, as the sphere is an unbounded but finite domain. The eigenvalue problem has no exact solution in the general case, however it has been extensively investigated through a variety of analytic approaches (e.g. [30, 31, 32, 33, 34, 35, 36]), and the eigenfrequencies and eigenfunctions can be well-approximated with analytical expressions in the different asymptotic regimes. In particular, in the regime $L_{\rm eq}\ll R$ (hereafter referred to as Matsuno limit), the spectrum of equatorial shallow-water waves exhibits two branches of modes with eastward group velocity (the equatorial Kelvin and Yanai waves) transiting through the frequency gap between Rossby waves and inertia-gravity waves. In contrast, those seem to be absent from the frequency spectrum in the other limit (see Figure 1), i.e. when $\Omega$ is small compared to $c/R$, hereafter referred to as Margules limit [30]. One would think that these branches somehow disappear as $L_{\rm eq}$ and $R$ become comparable, however there is no analytical solution of the eigenvalue problem in this intermediate situation to show exactly how.
@@
Figure 1: Dimensionless frequencies of the shallow-water model on the rotating sphere for two values of the parameter $\epsilon=c/\Omega R$ (left: $\epsilon=0.01$, strong rotation; right: $\epsilon=10$, weak rotation), in function of the azimuthal wave number $m\in\mathbb{Z}$, computed with Dedalus [37] (see Appendix 6.1 for numerical methods). In the Matsuno limit (left panel, $\epsilon\ll 1$), two distinct branches of frequencies transit through the gap between Rossby and inertia-gravity waves, whereas in the Margules limit (right panel, $\epsilon\gg 1$) these are clearly separated by an unfilled frequency gap.
@@
This paper intends to shed a new light on the loss of gapless Kelvin and Yanai modes, not with exact analytical resolution but by combining the study of the structure of modes with spectral topology. This analysis relies on numerous previous works in which topology brought new insights on wave properties in geophysical and astrophysical media [1, 18, 19, 20, 21, 38]. In the first part, we introduce the equations of the rotating shallow-water model on the sphere, and the relevant dimensionless parameters that characterise it. We recall the asymptotic solutions in both Matsuno and Margules limits. In the second part, we discuss the concept of modal flow and show that the latter is equal to zero for the spectrum of shallow-water waves on the rotating sphere, whereas it is equal to $+2$ on the unbounded $\beta$-plane. In the third part, we provide a topological interpretation of this result. We compute the Chern numbers that characterise the representation of the wave equations in a dual parameter space, which are associated to the points where the Weyl symbol of the wave operator has multiple eigenvalues. We show that the metric term induces degeneracy points of non-zero Chern numbers, in addition to the one exhibited in [1]. This analysis reconciles the bulk-interface correspondence established by Delplace et al on the unbounded $\beta$-plane [1] with the disappearance of the equatorial spectral flow in spherical geometry.
@@ 2.1 Linearised equations
We consider the linearised shallow-water equations of an inviscid fluid layer on top of a sphere of radius $R$, rigidly rotating with constant rate $\mathbf{\Omega}$ (Figure 2). The fluid is initially at rest in the rotating frame. Besides, assuming that the gravity $g$ at the surface is such that $g\gg R\Omega^{2}$, we ignore the centrifugal effects. Therefore, the rest state is that of a quiet layer of constant depth $h_{0}$. Noting the perturbed fields with ${}^{\prime}$, we define $\mathbf{v}=\mathbf{v}^{\prime}(\mathbf{x},t)$ and $h=h_{0}+h^{\prime}(\mathbf{x},t)$ the two-component velocity and height of the fluid, respectively, both functions of time $t$ and position $\mathbf{x}$ on the sphere. The linearised shallow-water equations can be conveniently expressed in terms of rescaled perturbation fields, i.e. $\tilde{\mathbf{v}}=\sqrt{h_{0}}\ \mathbf{v}^{\prime}=\tilde{u}\,\mathbf{e}_{\phi}+\tilde{v}\,\mathbf{e}_{\theta}$ and $\tilde{h}=\sqrt{g}\ h^{\prime}$, as
@@
where $c=\sqrt{gh_{0}}$ is the constant phase speed and $\mathbf{f}=(2\mathbf{\Omega}\cdot\mathbf{e}_{z})\mathbf{e}_{z}$ is the projection of $2\mathbf{\Omega}$ on the local unit vector $\mathbf{e}_{z}$ normal to the surface (traditional Coriolis parameter). Since none of the parameters appearing in Equations (2.1) depend on time or longitude, the rotating sphere acts as a waveguide trapping zonally-propagating waves in the meridional direction. We can thus expand any perturbation on the set of Fourier modes $X(\theta){\rm e}^{{\rm i}(m\phi-\omega t)}$, where $X$ is any of the three dependent variables of Equations (2.1), $(\phi,\theta)$ the longitudinal and latitudinal coordinates, respectively, and $(m,\omega)$ the azimuthal wave number and frequency associated with the Fourier mode, which corresponds to the solutions of
@@ 2.2 Lamb parameter and asymptotic solutions
If one rather considers the dimensionless frequency $\omega/\Omega$, the eigenvalue problem (2.1) depends only on one dimensionless parameter:
@@
The quantity $2/\epsilon$ is sometimes called the Lamb parameter [34], not to be confused with the Lamb frequency used in asteroseismology [39]. $1/\epsilon$ gives the traversal time of a gravity wave over the spherical body in planetary days. This parameter naturally appears in asymptotic expansions to compute the approximate spectrum of shallow-water waves [34, 36] or their ray paths [40] on the sphere. One can expect to asymptotically recover the Matsuno spectrum as $\epsilon\rightarrow 0$. Indeed, Expression (2) yields $L_{\rm eq}/R=\sqrt{\epsilon/2}$ for the equatorial radius of deformation, which means that, for small $\epsilon$, the eigenfunctions are equatorially trapped with angular spreading of order $\sqrt{\epsilon}$, thus recovering the Cartesian $\beta$-plane approximation made by Matsuno. In this limit, the solutions of (2.1) are appropriately described with Hermite polynomials, which are exact solutions on the unbounded $\beta$-plane (see 3.1).
@@
  • The ones such that $\omega=\mathcal{O}(c/R)$. One recovers the non-rotating shallow-water equations from (2.1). The eigenfunctions are those of the Laplacian operator on a prolate spheroid, i.e. the prolate spheroidal wave functions [6], and we have $R\omega/c\simeq\sqrt{\ell(\ell+1)}-m/\epsilon$ (see e.g. [41, 34]) with non-zero integers $\ell$ and $|m|\leq\ell$. These are global surface gravity oscillations on the sphere, perturbed by a small Coriolis term. For these solutions, the number of zeros of $\tilde{v}$ on the open interval $(-\pi/2,\pi/2)$ is given by $p=\ell-|m|+1$ [34].
  • The ones such that $\omega=\mathcal{O}(\Omega)$. These are quasi-geostrophic planetary (Rossby) modes. From (2.1), it is straightforward to check that they obey $\nabla\cdot\mathbf{v}^{\prime}=0$. Eliminating the other variables from Equations (4a), we end up with a stationary Schrödinger equation for $V=\cos(\theta)^{\frac{3}{2}}\,\tilde{v}$:
@@
    Solutions of Equation (6) are provided for instance in [42]. They yield
@@
    where the index $p$ indicates the number of zeros of the meridional velocity $\tilde{v}$ in the open interval $(-\pi/2,\pi/2)$. In the Margules limit, the largest Rossby frequency is thus $\omega=\Omega$, for $m=-1$ and $p=0$. The Rossby modes with $p=0$ constitute the westward Yanai modes [43]. Expression (7), which is exact for any azimuthal wave number $m$ in the Margules limit $\epsilon\gg 1$, is the same as (3.2c) obtained in [34] through a covariant formulation of the wave equations (2.1). Figure 3 provides a comparison of Expression (7) with numerical simulations, demonstrating the convergence of the Rossby waveband toward the Margules limit as $\epsilon$ increases.
@@
Figure 3: Comparison between Expression (7) (black triangles) and numerically-computed Rossby eigenfrequencies for $\epsilon=5$. The red dots are obtained by projecting Equations (2.1) on Dedalus’ curvilinear spectral basis [37], whereas the blue dots are computed from the problem (10) on $(-\pi/2,\pi/2)$, not necessarily with integer values of $m$, projecting the eigenfunctions on the basis of Chebyshev polynomials [44] (see Appendix 6.1). The accuracy between the numerically-converged frequencies for $\epsilon=5$ and Margules’ asymptotic formula (7) is approximately $0.4\%$.
@@ 2.4 Opening of a spectral gap as $\boldsymbol{\epsilon}$ increases
In the Margules limit ($\epsilon\gg 1$), the minimal inertia-gravity frequency is equal to $\sqrt{2}\>c/R$ and the maximum Rossby eigenfrequency is $\Omega$, which yields a frequency gap of width $\Omega\left(\sqrt{2}\>\epsilon-1\right)$ in which there is no mode. Conversely, in the Matsuno limit ($\epsilon\ll 1$), the frequency gap between Rossby and inertia-gravity modes is filled with the Yanai and Kelvin modes. However, talking about a frequency gap and branches of modes is abusive in this situation. Indeed, contrary to the zonal wave number $k$ of zonally-propagating waves on the unbounded $\beta$-plane, the azimuthal wave number $m$ is not a continuous parameter, since it only takes discrete integer values. Therefore the spectrum $(m,\omega)$ does not consist in continuous branches but rather a discrete set of points, which means that, strictly speaking, there are frequency gaps between any allowed values of $\omega$. Nevertheless, the equatorial Yanai and Kelvin modes define a clear connection between the Rossby and inertia-gravity wavebands on the unbounded $\beta$-plane, which is still visible in the shallow-water spectrum on the sphere with small $\epsilon$, even if $m$ takes discrete values (see Figures 1 and 4). Such a connection in the spectrum is referred to as spectral flow, which is usually defined for a continuous spectral parameter such as $k$ in unbounded flow models [1, 26, 18, 22, 19, 20, 21, 58]. The purpose of the following section is to extend the concept of spectral flow to the present situation, i.e. with a discrete wave number $m$, and show that the net number of eigenmodes gained by the shallow-water wavebands, when those modes are naturally labelled by the zeros of their meridional velocity, is actually zero for any value of $\epsilon$, even if there are transiting eigenfrequencies for small $\epsilon$ (see Figure 4).
@@
Figure 4: Numerically-computed eigenfrequencies $\omega$ for increasing values of $\epsilon$ and $|m|\leq 20$. These converged values are obtained by projecting Equations (2.1) on a basis of spin-weighted harmonics, a spectral method implemented in the Dedalus solver [37] (see Appendix 6.1). While for small $\epsilon$ two distinguishable spectral branches, the Yanai and Kelvin modes, transit through the gap between Rossby and gravity modes, these branches progressively break down and a clear gap opens up, separating Rossby and gravity frequencies.
3 Modal flow and meridional zeros of eigenfunctions

### v2

@@ Abstract
1Univ Lyon, ENS de Lyon, Univ Claude Bernard, CNRS, Centre de Recherche Astrophysique de Lyon (UMR CNRS 5574), F-69230 Saint-Genis-Laval, France
2Univ Lyon, ENS de Lyon, Univ Claude Bernard, CNRS, Laboratoire de Physique (UMR CNRS 5672), F-69342 Lyon, France
@@ 1 Introduction
However, the equatorial $\beta$-plane approximation used by [13, 1] (and many other studies of geophysical fluid dynamics) is limited in the sense that it only accounts for the curvature of the surface in the variation of the Coriolis parameter, at linear order, and not in the metric. It thus ignores the existence of the poles, and generates a spectrum of solutions defined on the unbounded domain across the equator, which seems paradoxical. Nevertheless, the solutions are accurate as long as they remain trapped around the equator, in the so-called Yoshida waveguide [14, 13], where the linear approximation holds. The trapping length, called the equatorial radius of deformation, is given by
@@
where $c$ is the speed of surface waves without rotation. $L_{\rm eq}$ thus needs to be much smaller than $R$, the radius of the planet. This condition is satisfied for fast-rotating planets, compared to the timescale of propagation of the waves at their surface. For the first baroclinic mode (i.e. the fastest one) in the equatorial ocean on Earth, the equatorial radius of deformation is approximately $300$ kilometers (see e.g. [3], p. 304), so the $\beta$-plane approximation is quite accurate in this context. However, when it comes to global oscillations of larger scales, both the curved metric and the actual sine variation of $f$ with latitude must be taken into account. Moreover, the quantisation of the zonal wave number $k$ cannot be ignored, especially when the wavelength is not small compared to the radius $R$. The full spectrum of shallow-water waves on the sphere is a peculiar problem, as the sphere is an unbounded but finite domain. The eigenvalue problem has no exact solution in the general case, however it has been extensively investigated through a variety of analytic approaches (e.g. [15, 16, 17, 18, 19, 20, 21]), and the frequencies and wave functions can be well-approximated with analytical expressions in the different asymptotic regimes. In particular, in the regime $L_{\rm eq}\ll R$ (hereafter referred to as Matsuno limit), the spectrum of equatorial shallow-water waves is well-approximated by Matsuno’s spectrum and thus clearly exhibits two branches of modes with eastward group velocity transiting through the frequency gap between the different wavebands (in the rest of the paper, we will refer to the different groups of shallow-water modes as wavebands. There are three wavebands in the shallow-water model: the Rossby waveband and two equivalent inertia-gravity wavebands). In contrast, those seem to be absent from the frequency spectrum in the other limit (see Figure 1), i.e. when $\Omega$ is small compared to $c/R$, hereafter referred to as Margules limit [15]. One would think that these branches somehow disappear as $L_{\rm eq}$ and $R$ become comparable, however there is no analytical solution of the eigenvalue problem in this intermediate situation to show exactly how.
@@
Figure 1: Dimensionless frequencies of the shallow-water model on the rotating sphere for two values of the parameter $\epsilon=c/\Omega R$ (left: $\epsilon=0.01$, strong rotation; right: $\epsilon=10$, weak rotation), in function of the azimuthal wave number $m\in\mathbb{Z}$, computed with Dedalus [22] (see Appendix 6.1 for numerical methods). In the Matsuno limit (left panel, $\epsilon\ll 1$), two distinct branches of frequencies transit through the gap between Rossby and inertia-gravity waves, whereas in the Margules limit (right panel, $\epsilon\gg 1$) these are clearly separated by an unfilled frequency gap. Note that the frequency axes in this figure are in logarithmic scale, as well as in Figure 4, in order to visually appreciate the transition between the Matsuno and Margules regimes. All the other frequency plots in the article are displayed in linear scale.
Alternatively, a few years ago, the presence of these two transiting branches in the spectrum of shallow-water waves on the unbounded $\beta$-plane was interpreted as a manifestation of an underlying topological property. Indeed, after noticing the strong resemblance of this $+2$ spectral flow of modes with the topologically-protected modes crossing the gap in certain insulating materials, [1] applied the same arguments used in condensed matter physics to define topological integers associated to this spectral flow, the Chern numbers or topological charges. These integers characterise the phase singularities of Kelvin’s plane-wave solutions, which are computed for constant $f$ and thus require only to diagonalise a 3-by-3 matrix. Nevertheless, they happen to be equal to the number of modes gained by the associated wavebands in the more elaborated $\beta$-plane model. This counter-intuitive result, which connects the topological properties of Kelvin’s $f$-plane modes to a spectral property of Matsuno’s $\beta$-plane spectrum, is a consequence of the more general index theorem [23, 24, 25]. It has been applied to predict the conditions of existence and number of modes transiting across the frequency gap in a variety models from a great variety of domains [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
@@
This paper intends to shed a new light on the loss of gapless Kelvin and Yanai modes on the sphere, not with exact analytical resolution but by combining the study of the structure of modes with spectral topology. This analysis relies on numerous previous works in which topology brought new insights on wave properties in geophysical and astrophysical media [1, 26, 27, 28, 29, 37]. In the first part, we introduce the equations of the rotating shallow-water model on the sphere, and the relevant dimensionless parameters that characterise it. We recall the asymptotic solutions in both Matsuno and Margules limits. In the second part, we discuss the concept of modal flow and show that the latter is equal to zero for the spectrum of shallow-water waves on the rotating sphere, whereas it is equal to $+2$ on the unbounded $\beta$-plane. In the third part, we provide a topological interpretation of this result. We compute the Chern numbers, which are associated to the points where the symbol of the wave operator has multiple eigenvalues. We show that the metric term induces degeneracy points of non-zero Chern numbers, in addition to the one exhibited by [1]. This analysis reconciles the bulk-interface correspondence established by Delplace et al on the unbounded $\beta$-plane with the disappearance of the equatorial spectral flow in spherical geometry.
@@ 2.1 Linearised equations
We consider the linearised shallow-water equations of an inviscid fluid layer on top of a sphere of radius $R$, rigidly rotating with constant rate $\mathbf{\Omega}$ (Figure 2). The fluid is initially at rest in the rotating frame. Besides, assuming that the gravity $g$ at the surface is such that $g\gg R\Omega^{2}$, we ignore the centrifugal effects. Therefore, the rest state is that of a quiet layer of constant depth $h_{0}$. Noting the perturbed fields with , we define $\mathbf{v}=\mathbf{v}^{\prime}(\mathbf{x},t)$ and $h=h_{0}+h^{\prime}(\mathbf{x},t)$ the two-component velocity and height of the fluid, respectively, both functions of time $t$ and position $\mathbf{x}$ on the sphere. The linearised shallow-water equations can be conveniently expressed in terms of rescaled perturbation fields, i.e. $\tilde{\mathbf{v}}=\sqrt{h_{0}}\ \mathbf{v}^{\prime}=\tilde{u}\,\mathbf{e}_{\phi}+\tilde{v}\,\mathbf{e}_{\theta}$ and $\tilde{h}=\sqrt{g}\ h^{\prime}$, as
@@
where $c=\sqrt{gh_{0}}$ is the constant phase speed and $\mathbf{f}=(2\mathbf{\Omega}\cdot\mathbf{e}_{z})\mathbf{e}_{z}$ is the projection of $2\mathbf{\Omega}$ on the local unit vector $\mathbf{e}_{z}$ normal to the surface (traditional Coriolis parameter). Since none of the parameters appearing in Equations (3) depend on time or longitude, the rotating sphere acts as a waveguide trapping zonally-propagating waves in the meridional direction. We can thus expand any perturbation on the set of Fourier modes $X(\theta){\rm e}^{{\rm i}(m\phi-\omega t)}$, where $X$ is any of the three dependent variables of Equations (3), $(\phi,\theta)$ the longitudinal and latitudinal coordinates, respectively, and $(m,\omega)$ the azimuthal wave number and frequency associated with the Fourier mode, which corresponds to the solutions of
@@ 2.2 Lamb parameter and asymptotic solutions
If one rather considers the dimensionless frequency $\omega/\Omega$, the eigenvalue problem (4) depends only on one dimensionless parameter:
@@
The quantity $2/\epsilon$ is sometimes called the Lamb parameter [19], not to be confused with the Lamb frequency used in asteroseismology [38]. $1/\epsilon$ gives the traversal time of a gravity wave over the spherical body in planetary days. Equations (4) in dimensionless units are
$(\omega/\Omega)\tilde{\mathbf{v}}=-{\rm i}\epsilon\tilde{\nabla}\tilde{h}-2{\rm i}\sin(\theta)\mathbf{e}_{z}\times\tilde{\mathbf{v}}\ ,$ (6a)
$(\omega/\Omega)\tilde{h}=-{\rm i}\epsilon\tilde{\nabla}\cdot\tilde{\mathbf{v}}\ ,$ (6b)
where ~$\tilde{\nabla}=R\nabla$ is the gradient operator on the unit sphere. The parameter $\epsilon$ naturally appears in asymptotic expansions to compute the approximate spectrum of shallow-water waves [19, 21] or their ray paths [39] on the sphere. One can expect to asymptotically recover the Matsuno spectrum as $\epsilon\rightarrow 0$. Indeed, Expression (2) yields $L_{\rm eq}/R=\sqrt{\epsilon/2}$ for the equatorial radius of deformation, which means that, for small $\epsilon$, the wave functions are equatorially trapped with angular spreading of order $\sqrt{\epsilon}$, thus recovering the Cartesian $\beta$-plane approximation made by Matsuno. In this limit, the solutions of (4) are appropriately described with Hermite polynomials, which are exact solutions on the unbounded $\beta$-plane (see 3.1).
@@
  • The ones such that $\omega=\mathcal{O}(c/R)$. One recovers the non-rotating shallow-water equations from (4). The wave functions are those of the Laplacian operator on a prolate spheroid, i.e. the prolate spheroidal wave functions [6], and we have $R\omega/c\simeq\sqrt{\ell(\ell+1)}-m/\epsilon$ (see e.g. [40, 19]) with non-zero integers $\ell$ and $|m|\leq\ell$. These are global surface gravity oscillations on the sphere, perturbed by a small Coriolis term. For these solutions, the number of zeros of $\tilde{v}$ on the open interval $(-\pi/2,\pi/2)$ is given by $p=\ell-|m|+1$ [19].
  • The ones such that $\omega=\mathcal{O}(\Omega)$. These are quasi-geostrophic planetary (Rossby) modes. From (4), it is straightforward to check that they obey |~$|\tilde{\nabla}\cdot\mathbf{v}^{\prime}|=\mathcal{O}\left(|\mathbf{v}^{\prime}|/\epsilon^{2}\right)$, i.e. that these modes have asymptotically divergence-free velocity. Eliminating the other variables from Equations (4a), we end up with a second-order differential equation for $V=\cos^{\frac{3}{2}}(\theta)\,\tilde{v}$:
@@
    Solutions of Equation (7) are provided for instance by [41]. They yield
@@
    where the index $p$ indicates the number of zeros of the meridional velocity $\tilde{v}$ in the open interval $(-\pi/2,\pi/2)$. In the Margules limit, the largest Rossby frequency is thus $\omega=\Omega$, for $m=-1$ and $p=0$. The Rossby modes with $p=0$ constitute the westward Yanai modes [42]. Expression (8), which is exact for any azimuthal wave number $m$ in the Margules limit $\epsilon\gg 1$, is the same as (3.2c) obtained by [19] through a covariant formulation of the wave equations (3). Figure 3 provides a comparison of Expression (8) with numerical simulations, demonstrating the convergence of the Rossby waveband toward the Margules limit as $\epsilon$ increases.
@@
Figure 3: Comparison between Expression (8) (black triangles) and numerically-computed Rossby frequencies for $\epsilon=5$. The red dots are obtained by projecting Equations (4) on Dedalus’ curvilinear spectral basis [22], whereas the blue dots are computed from the problem (11) on $(-\pi/2,\pi/2)$, not necessarily with integer values of $m$, projecting the wave functions on the basis of Chebyshev polynomials [43] (see Appendix 6.1). The accuracy between the numerically-converged frequencies for $\epsilon=5$ and Margules’ asymptotic formula (8) is approximately $0.4\%$.
@@ 2.4 Opening of a spectral gap as $\boldsymbol{\epsilon}$ increases
In the Margules limit ($\epsilon\gg 1$), the minimal inertia-gravity frequency is equal to $\sqrt{2}\>c/R$ and the maximum Rossby frequency is $\Omega$, which yields a frequency gap of width $\Omega\left(\sqrt{2}\>\epsilon-1\right)$ in which there is no mode. Conversely, in the Matsuno limit ($\epsilon\ll 1$), the frequency gap between Rossby and inertia-gravity modes is filled with the Yanai and Kelvin modes. However, talking about a frequency gap and branches of modes is abusive in this situation. Indeed, contrary to the zonal wave number $k$ of zonally-propagating waves on the unbounded $\beta$-plane, the azimuthal wave number $m$ is not a continuous parameter, since it only takes discrete integer values. Therefore the spectrum $(m,\omega)$ does not consist of continuous branches but rather a discrete set of points, which means that, strictly speaking, there are frequency gaps between any allowed values of $\omega$. Nevertheless, the equatorial Yanai and Kelvin modes define a clear connection between the Rossby and inertia-gravity wavebands on the unbounded $\beta$-plane, which is still visible in the shallow-water spectrum on the sphere with small $\epsilon$, even if $m$ takes discrete values (see Figures 1 and 4). Such a connection in the spectrum is referred to as spectral flow, which is usually defined for a continuous spectral parameter such as $k$ in unbounded flow models [1, 34, 26, 30, 27, 28, 29, 57]. The purpose of the following section is to extend the concept of spectral flow to the present situation, i.e. with a discrete wave number $m$, and show that the net number of modes gained by the inertia-gravity waveband, when those modes are naturally labelled by the zeros of their meridional velocity, is actually zero for any value of $\epsilon$, even if there are transiting frequencies for small $\epsilon$ (see Figure 4).
@@
Figure 4: Numerically-computed frequencies $\omega$ for increasing values of $\epsilon$ and $|m|\leq 20$. These converged values are obtained by projecting Equations (4) on a basis of spin-weighted harmonics, a spectral method implemented in the Dedalus solver [22] (see Appendix 6.1). While for small $\epsilon$ two distinguishable spectral branches, the Yanai and Kelvin modes, transit through the gap between Rossby and gravity modes, these branches progressively break down and a clear gap opens up, separating Rossby and gravity frequencies.
3 Modal flow and zeros of the meridional velocity